OpenStudy (anonymous):

Three points A , B , C are on a straight line such Ab = BC . A particle moving acceleration passes A and C with speeds of 6m/s and 8 m/s. Another particle moving with uniform acceleration passes A and B with speed of 12m/s and 14m/s if t1 and t2 are time taken by the particle to cross AB t1/t2 is nearly?

OpenStudy (anonymous):

The answer should be 2

OpenStudy (ajprincess):

time taken by the particles or particle?

OpenStudy (anonymous):

It is not given we have to find it

OpenStudy (anonymous):

Can anyone answer This!

OpenStudy (anonymous):

@ajprincess can uanswer this

OpenStudy (anonymous):

yeah, the answer is correct

OpenStudy (anonymous):

Whuch answer?

OpenStudy (anonymous):

the answer should be 2^^

OpenStudy (anonymous):

u know the basic kinematics formulae?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, say, the time taken are t1 and t2 and accelerations are a1, a2 respectively

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now, apply v^2=u^2+2as

OpenStudy (anonymous):

as s is fixed, u can get the ratio of their acceleration

OpenStudy (anonymous):

sorry, s is 2 times in one case^^

OpenStudy (anonymous):

can u get the ratio?

OpenStudy (anonymous):

I gave Up !! NO Idea

OpenStudy (anonymous):

i am disappointed.. u r not trying

OpenStudy (anonymous):

i will try my best

OpenStudy (anonymous):

say, the first person's acceleration and time are a1, t1 second person's acceleration and time are a2, t2 say, AB=s then applying the kinematics formulae.. 8^2=6^2+2a1*2s or s=7/a1 14^2=12^2+2a2*s or s=26/a2 now, equate s from both cases 7/a1=26/a2 a2/a1=26/7 got it so far?

OpenStudy (anonymous):

i got a2=13a/7

OpenStudy (anonymous):

no, look carefully, first person covers AC, second person covers AB so, first person goes 2s, second one goes s

OpenStudy (anonymous):

in question it is said that AB = Bc

OpenStudy (anonymous):

@Arnab09 so diatances are same

OpenStudy (anonymous):

so, AC=2AB

OpenStudy (anonymous):

yeah u r correct!

OpenStudy (anonymous):

Plz go on !

OpenStudy (anonymous):

okay, then apply: v=u+at so, 8=6+a1*t1 t1=2/a1 14=12+a2*t2 so, t2= 2/a2 so, t1/t2= a2/a1 as proved before, a2/a1= 26/7 so, t1/t2= 26/7 answer will be 4 approx got it?

OpenStudy (anonymous):

@Arnab09 the answer should be 2

OpenStudy (anonymous):

@mahmit2012 i think u r correct

OpenStudy (anonymous):